Express the roots of the equation z3 −α3 = 0 in terms of α and w, where w is a complex cube root of unity. Use your answer to find the roots of the following equations in the form a + ib.
1
Expert's answer
2020-05-19T09:13:06-0400
"z^3-\alpha^3=0"

"z^3-\alpha^3=(z-\alpha)(z^2+\alpha z+\alpha^2)"

Then


"(z-\alpha)(z^2+\alpha z+\alpha^2)=0"

"z_1=\alpha\" or "z^2+\alpha z+\alpha^2=0"


"z={-\alpha\pm\sqrt{\alpha^2-4\alpha^2}\over2}=\alpha({-1\pm j\sqrt{3}\over2})"

Cube Root of Unity Value


"w_1=1,\ real""w_2={-1- j\sqrt{3}\over2}, complex""w_3={-1+ j\sqrt{3}\over2}, \ complex"

"z_1=\alpha w_1=\alpha\cdot1=1\cdot\alpha+j\cdot0"

"z_2=\alpha w_2=\alpha\cdot{-1- j\sqrt{3}\over2}=-{1\over 2}\alpha-j\cdot{\alpha\sqrt{3}\over 2}"

"z_2=\alpha w_3=\alpha\cdot{-1+ j\sqrt{3}\over2}=-{1\over 2}\alpha+j\cdot{\alpha\sqrt{3}\over 2}"